### v1

# Non-iterative triples for transcorrelated coupled cluster

Abstract
We present an implementation of a perturbative triples correction for coupled cluster including single and double excitations based on the transcorrelated Hamiltonian. Transcorrelation introduces explicit electron correlation in the electronic Hamiltonian through similarity transformation with a correlation factor. Due to this transformation, the transcorrelated Hamiltonian includes up to three-body couplings and becomes non-hermitian. Since conventional coupled cluster is solved by projection, it is well suited to harbor non-hermitian Hamiltonians. The arising three-body operator, however, creates a huge memory bottleneck and increases the runtime scaling of the coupled cluster equations. As it has been shown that the three-body operator can be approximated, by expressing the Hamiltonian in the normal-ordered form, we investigate this approximation for the perturbative triples correction. Results are compared with code-generation based transcorrelated coupled cluster up to quadruple excitations.
1 Introduction
The accurate description of electron correlation remains one of the greatest challenges in quantum chemistry. Within a one-particle picture, the energy contribution of electron correlation can be defined as the difference between the exact full configuration interaction (FCI), and the mean-field Hartree–Fock (HF) energy. One of the most powerful approaches to recover the correlation energy is the coupled cluster1, 2, 3, 4 (CC) approach. In CC, the wave function is parameterized by an exponential cluster operator acting on a reference wave function (usually the HF determinant). In practice, the respective wave function expansion is truncated after single and double excitations5, 6 (CCSD) to avoid the exponential scaling. The inclusion of full triple excitations (CCSDT)7, 8 can significantly improve on the accuracy of the correlation energy9, but accounting for it is computationally too demanding for most system sizes. Hence, triples are usually approximated by perturbative approaches10 (in particular in the CCSD(T) model) as a good compromise between cost and accuracy.
The size of the one-particle basis, given by a set of Gaussian functions, determines the overall quality of the a calculation. With an increasing number of Gaussians, the basis set eventually converges to the complete basis set limit. However, due to the high computational scaling of most ab initio methods, it is not always feasible to apply large basis sets.
Because it enforces the electron-electron cusp 11, 12, the explicit inclusion of the electronic distance in the wave function can accelerate the basis set convergence11, therefore increasing the accuracy of calculations with a given basis set. For most applications, the Ansätze of the / family13, 14, 15 provide a way to explicitly correlate electrons within the wave function. However, a major drawback of these methods is the non-trivial generalization to higher than doubles excitations. In most implementations, the perturbative triples correction to CCSD- is evaluated by the standard CCSD(T) formula (see Eq. (25) below), which leads to the finite basis set error dominating other potential errors. Werner and coworkers addressed this issue, by scaling the perturbative triples correction by the basis set correction of Møller-Plesset pertubation theory, which assumes the same convergence rate to the complete basis set limit and breaks the size-consistency16, 17. The size-consistency issue was later addressed by Kallay et al.18. A different approach exists, which introduced an explicitly correlated component into the expression for the triples correction19, 20. However, this approach did not yield improvements comparable to CCSD-.
A different approach to explicitly correlate electrons is transcorrelation, introduced by Boys and Handy21, which we adopt in this study. In contrast to including the inter-electronic distance into the wave function, the Hamiltonian is similarity transformed by a correlation factor. As a result, transcorrelation is generally applicable, even to multi-configurational wave functions 22, 23, 24, 25. However, the resulting transcorrelated Hamiltonian is no longer hermitian and includes a three-body operator. Recently, the transcorrelation idea has experienced a vivid revival26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 24, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 25, 59, 60, 61, 62, 63.
In the context of CC, the non-hermiticity does not impose any issues, since the Hamiltonian is similarity transformed by the exponential cluster operator, resulting already in a non-hermitian Hamiltonian. Nevertheless, the full treatment of the three-body operator increases the computational cost and memory scaling. Ten-no combined the transcorrelated Hamiltonian with linearized CC64, where many contractions involving the three-body operator were neglected. Later Alavi and coworkers applied CCSD to the transcorrelated Hamiltonian (TC-CCSD) for the homogeneous electron gas49, atoms50 and molecules51. In order to deal with the three-body couplings, they introduced an approximation based on the normal-ordered Hamiltonian51. In this approximation, the full TC Hamiltonian is normal-ordered with respect to a single reference determinant and subsequently the three-body integrals are neglected. Within this procedure, mean-field contributions are transferred from the three-body operator to two-, one-, and zero-body operators. In the same paper, they introduced a similar approximation based on the -dressed Hamiltonian65, 66, which, however, still suffers from the huge memory scaling. In the -dressed Hamiltonian approximation, all remaining three-body terms are neglected, meaning that effective intermediates are pre-contracted with the three-body operator before the operator is neglected. Another approach67 to reduce the drawbacks of the three-body operator is based on the generalized normal-ordering68, 69 of the three-body operator, which reduces the scaling to .
For higher-order CC models, the explicit treatment of the three-body operator results in numerous possible diagrams. A derivation of the CCSD equation for Hamiltonians including three-body couplings was given by Piecuch and coworkes in the context of nuclear structure theory70. The total number of CCSD diagrams for a Hamiltonian including three-body couplings increases from 48 to 116. Hence, tools for automatic code contractions become important for the implementation of TC-CC with higher than doubles excitations.
In this work, we derive a non-iterative triples approximation for TC-CCSD(T) including three-body operators and the normal-order approximation. A first CCSD(T) implementation for Hamiltonians including a three-body operator71 was derived based on the CCSD(T)72, 73, 74 method. Later, Kats et al. derived the CCSD(T) method for the transcorrelated Hamiltonian62, however without the inclusion of the full three-body operator.
We implemented TC-CCSDT and TC-CCSDTQ for the full transcorrelated Hamiltonian with the Wick&d75, 76 library to generate the required CC tensor contractions. In the following section, we first briefly review the conventional non-iterative triples approximation and the transcorrelated Hamiltonian. Afterwards, we show the required diagrams including three-body couplings for perturbative triples. The correlation factor applied in this work is introduced in section 3. We compare these TC-CC models for Be, LiH and Be2 in section 4.
2 Theory
If not stated otherwise, we assume spin-orbital labels for all operators. For convenience, we summarize our notation in Table 1.
| Symbol | Meaning |
| occupied spin-orbital indices | |
| virtual spin-orbital indices | |
| general spin-orbital indices | |
| denotes a normal-ordered string of second quantized operators | |
| electron repulsion integrals in physics notation | |
| anti-symmetrized two-body integrals, e.g. |
The non-relativistic normal-ordered Hamiltonian in spin-orbital notation reads
where
is the mean-field energy contribution of the reference determinant, including the nuclear repulsion energy . is an element of the Fock matrix
where comprises all one-body operators, i.e. the kinetic energy and electron-nuclear interactions. In order to distinguish from the transcorrelated Hamiltonian, the former quantity is referred to as the conventional Hamiltonian. We consistently map the indices of the creation and annihilation operators by placing them in the superscript and subscript, respectively, in the corresponding tensor (bra and ket of anti-symmetrized integrals, respectively).
We employ the usual diagrammatic representation77, 78, 79, where dashed lines represent operators. For example, two outgoing arcs connected to an “x” correspond to a one-body operator, two outgoing arcs connected to another set of two outgoing arcs represent a two-body operator. Solid horizontal lines represent amplitudes, where the number of vertices corresponds to the excitation degree of the amplitude. The dotted line represents the denominator constructed from the diagonal entries of the Fock operator; for example
is the three-body denominator. The coupled cluster amplitudes are denoted by the usual symbols , , , and so forth. For convenience, we will employ both an algebraic and diagramatic representation in this work.
2.1 Conventional perturbative triples
In this section, we re-derive the conventional perturbative triples correction to the energy in the formalism of Stanton80. In subsequent section, this framework is extended to incorporate the three-body operators in the electronic Hamiltonian.
As described by Stanton in Ref. 80, CCSD can be transformed into an eigenvalue problem of ,
where the eigenvector is the reference determinant . However, since the similarity transformation is not unitary, is not hermitian. Hence, the left eigenvector is not given by
In order to evaluate the left eigenvector explicitly, the so-called lambda equations have to be solved, so that
with where is a de-excitation operator. Now, instead of , is the zeroth-order Hamiltonian with and as left- and right-hand states, respectively. The perturbative expansion of the similarity transformed Hamiltonian according to Stanton80 results in the leading-order correction to the energy
where S, D, and T correspond to singly, doubly, and triply excited determinants, respectively, and D3 is a diagonal matrix with the inverse of the elements defined in Eq. (4) on the diagonal. This perturbative expansion leads to two different sets of amplitudes which can be evaluated according to
To reduce the computational cost, is usually approximated by , leading to two intermediates
Note that due to the commutator in Eq. (11) only connected diagrams can be created, while in Eq. (12) also disconnected diagrams appear. The tensor contractions for the first and second intermediate have the following form
and
respectively. Note that the last term in the latter equation vanishes due to the Brillouin condition in the case of conventional Hamiltonian, but this is no longer true for the transcorrelated Hamiltonian considered in the subsequent sections. Because of the anti-symmetry of the two-body integrals and the approximation of Eqs. (20) and (24) are equivalent. Therefore, in practice, the latter two diagrams in Eq. (20) are added to Eq. (15), so that the first tensor contraction must be evaluated only once. The energy contribution of the non-iterative triples correction for the conventional Hamiltonian can be written as
The skeleton diagrams in Eq. (29) can be derived by contracting the two intermediates with each other.
2.2 Transcorrelated Hamiltonian
The TC Hamiltonian can be derived by similarity transformation of the conventional Hamiltonian with a Jastrow factor81 of the form
resulting in the additional two-body term and the three-body term . For the sake of completeness, we summarize some of the important aspects of the TC Hamiltonian. In the following, we redefine
Since the two-body term is not hermitian, the re-defined, anti-symmetrized two-body integrals, lack the symmetry
The anti-symmetrized from of the three-body tensor can be defined as
such that
After normal-ordering of the three-body term
the one- and two-body components can be subtracted from the corresponding one- and two-body operators. Hence, the normal-ordered TC Hamiltonian in the spin-orbital basis reads
where
with
and
include a mean-field contribution from the three-body operator.
To derive the energy and amplitude equations, is similarity transformed with the cluster operator.
Because of the new three-body operator, the resulting Baker-Campbell-Hausdorf-series truncates only after the six-fold commutator.
To tackle the memory requirements associated with the three-body operator, the transcorrelated Hamiltonian can be approximated by the normal-ordering approximation50,
2.3 Transcorrelated perturbative triples
By application of Eqs. (11) and (12), we can derive a new set of diagrams for the intermediates of the perturbative triples with the three-body operator
Since the contribution of the three-body operator does not result in any disconnected diagrams (such disconnected diagrams would be of an order higher than three) and since the three-body operator is hermitian, this contribution is identical for both sets of amplitudes ( and ). Adding the new set of diagrams to the intermediates in Eqs. (20) and (15) yields the new set of equations:
However, the lack of hermiticity of the two-body operator in Eq. (32) has some implications for the formulas. First, due to the normal-ordering of the Hamiltonian, the one-body operator is no longer symmetric
and second, the contractions in Eqs. (20) and (24) are not equivalent anymore. Note that in standard implementations of the (T) correction with the conventional Hamiltonian, only one set of amplitudes, namely , is evaluated explicitly. The additional terms present in are inexpensive and can be evaluated on the fly, such that is never explicitly formed. By contrast, in the implementation of the (T) correction in the transcorrelated framework, both sets of amplitudes ( and ) must be separately evaluated. However, we do not expect this to cause a serious memory bottleneck, because these amplitudes can be evaluated in batches with some indices fixed, similarly as in the conventional case.
3 Computational Methodology
As the CC method proposed in this work is generally applicable for any correlator, we chose the same correlation factor as in Ref. 23, namely
and set .
As shown in Ref. 23, the following basic quantities are necessary to evaluate the three-body component of the transcorrelated Hamiltonian
where we have switched to the chemists’ notation for brevity. Here and throughout this section, Greek letters denote indices corresponding to the atomic orbitals basis. To make the evaluation of feasible, it has been proposed 23 to employ a combination of density-fitting (DF) and resolution of identity (RI) techniques. This leads to the following working formula
where the intermediates read
In the above expressions the indices denote elements of the auxiliary DF basis, while stand for the elements of the RI basis. As an unfortunate side effect of the RI approximation, the three-body integrals evaluated according to Eq. 52 no longer strictly obey the same symmetry relations (with respect to exchange of the orbital indices) as the original integrals . While this problem vanishes in the limit of a complete RI basis, it leads to technical complications, such as increased storage requirements, for any finite RI basis. To fix this issue, we introduce the following symmetrized intermediates
and use the redefined quantities and in evaluation of Eq. 52. One can show by direct calculation that this approach leads to the correct symmetry of the approximate three-body integrals, i.e. identical to the original quantities .
For all calculations, we selected the cc-pVDZ, cc-pCVDZ, cc-pVTZ and cc-pVQZ basis 82, 83 with the matching cc-pV{D, T}Z-RIFIT basis 84, 85 for the integral evaluation. If not stated otherwise, we employ the cc-pV5Z-RIFIT basis set for Li 85 and the uncontracted cc-pV5Z basis for H and Be 82, 83 in the resolution of identity approximation in the integral evaluation.
To implement the amplitude equations in spin-orbital and spin-factorized form, we employ the open-source library Wick&d 75, 76. Because of the flexibility of this library, it was possible to generate CC equations for any excitation degree and Hamiltonian with -body operators. Since the underlying algorithm is based on an anti-symmetrized form of the tensors with a spin-orbital representation of the creation and annihilation operators, the stored Hamiltonian and amplitudes include many entries, which are zero due to spin. Especially in the context of higher-order amplitudes or Hamiltonians including a three-body operator, deriving pure spin-orbital form of the amplitude equations is unfeasible because of the the resulting memory and runtime requirements . Hence, we generated the spin-factorized form of the amplitude equations, by creating two fermionic spaces, i.e. and -spin and declaring the required operators accordingly, as described in Ref. 75. If TC-CC is approximated by normal-ordering (TC-CC-n), the Hamiltonian include up to two-body couplings and the diagrams become equal to conventional CC.
The number of generated tensor contractions per CC truncation level can be seen for both spin-orbital and spin-factorized forms in Table 2. Due to the large number of required contractions for spin-factorized conventional CCSDT and CCSDTQ, the number of diagrams is enormous for these methods if three-body operators are included. Hence, we employed the spin-factorized form for all methods, but TC-CCSDT and TC-CCSDTQ.
All CC energies were converged up to Hartree.
| Diagrams (conv.) | Diagrams (TC) | |||||||||
| CC truncation | E | E | ||||||||
| Spin-orbital basis | ||||||||||
| 2 | 3 | 14 | 31 | 5 | 29 | 82 | ||||
| 3 | 3 | 15 | 37 | 47 | 6 | 35 | 111 | 224 | ||
| 4 | 3 | 15 | 38 | 53 | 74 | 6 | 36 | 117 | 253 | 410 |
| Spin-factorized basis | ||||||||||
| 2 | 8 | 52 | 172 | 18 | 164 | 740 | ||||
| 3 | 8 | 58 | 222 | 520 | ||||||
| 4 | 8 | 58 | 231 | 592 | 1289 | |||||
4 Results
We first investigated the basis set convergence of CCSD, CCSD(T), CCSDT and CCSDTQ with the conventional and transcorrelated Hamiltonian for the Be atom. Second, we compared theses methods in different basis sets for the dissociation of the LiH molecule. Finally, we studied the Be2 dimer with CCSD, CCSD(T) and CCSDT, as Be2 is a prototypical system that requires at least perturbative triples and a triple- basis set to be described correctly.
4.1 Be atom
Since the Be atom is a four electron system, CCSDTQ yields FCI accuracy. We compared the basis set convergence with a highly accurate result from Ref. 86. The normal-ordering approximation of the three-body and the effect of the CC truncation for different basis sets are shown in Table 3.
| Basis | Hamiltonian | CCSD | CCSD(T) | CCSDT | CCSDTQ |
| cc-pVDZ | conv. | -14.617532 | -14.617570 | -14.617570 | -14.617572 |
| cc-pVTZ | conv. | -14.623581 | -14.623812 | -14.623823 | -14.623832 |
| cc-pVQZ | conv. | -14.639631 | -14.640128 | -14.640161 | -14.640169 |
| cc-pVDZ | TC-n | -14.656796 | -14.656787 | -14.656803 | -14.656806 |
| cc-pVDZ | TC | -14.656813 | -14.656805 | -14.656821 | -14.656823 |
| cc-pVTZ | TC-n | -14.658693 | -14.658747 | -14.658791 | -14.658797 |
| cc-pVTZ | TC | -14.658713 | -14.658767 | -14.658809 | -14.658816 |
| CBS FCI86 | -14.667356 | ||||
Table 3 shows the expected behavior of the total electronic energy with respect of CC truncation level and basis set size: the inclusion of perturbative triples improved the CCSD energies, with a further improvement by the full triples. The impact of the perturbative triples increased with basis set size, reducing the error of CCSD with respect to CCSDTQ from 0.04 mHa to 0.002 mHa, 0.251 mHa to 0.02 mHa, and 0.537 mHa to 0.041 mHa, for the double-, triple-, and quadruple- basis, respectively. In contrast to CCSD(T), CCSDT only yielded a negligibly small further improvement.
For the transcorrelated Hamiltonian in double- basis, the application of the perturbative triples slightly detoriated the CCSD energies with respect to CCSDTQ. However, the absolute error for CCSD is only 0.01 mHa and for CCSD(T) 0.019 mHa with respect to CCSDTQ. In the larger triple- basis, CCSD(T) improves on the CCSD energies, reducing the error from 0.103 mHa to 0.048 mHa and 104 mHa to 0.05 mHa for the full transcorrelated and normal-ordering approximated methods, respectively. Even for the same basis set size, CCSD captured more correlation energy with the transcorrelated Hamiltonian than with the conventional Hamiltonian. The normal-ordering approximation yielded results similar to those obtained with the full transcorrelated Hamiltonian, independent of the CC truncation level, while reducing the overall memory scaling.
4.2 LiH potential energy curve
We studied the LiH as a potential single-reference diatomic system. Hence, higher-order clusters should only have a small effect on the potential energy curve. Like the Be atom, LiH is a four electron system, hence CCSDTQ energies are of FCI accuracy.
We compare results obtatined with the transcorrelated with the conventional Hamiltonian for the different CC models in Fig. 1. Since the LiH dissociation can be reliably described by CCSD, higher-order cluster contributions, as in CCSD(T), CCSDT and CCSDTQ, do not improve the energy convergence significantly. The potential energy curves in Fig. 1 confirm this, because the various CC methods resulted in a similar potential energy curve progression. We also observed similar trends as those found for the transcorrelated density matrix renormalization group in Ref. 63. First, the results for the transcorrelated Hamiltonian improves the basis set convergence by up to two cardinal numbers. Second, the effect of the normal-ordering approximation did not significantly affect the energy, which confirms the results from previous work50, 63. And third, the transcorrelated potential energy curves are not parallel to the conventional curves, which might be due to our choice of the correlation factor. However, in contrast to other correlation factors, our choice is universal and simpler, which is an advantage for the definition of a universal electronic structure model. In view of our results with specific core-valence basis sets, we presume that the parallelity issue can be solved by fitting proper atomic orbital basis sets for a transcorrelated approach, rather than introducing parameters into the correlators that may then be optimized (see also our results for the Be dimer in the next section).
Next we investigate the dissociation of the Be2 dimer, to compare our perturbative triples approximation with respect to full triple excitations and to study the role of the basis set.
4.3 Be2 potential energy curve
A qualitatively correct description of the potential energy curve of the beryllium dimer with the conventional Hamiltonian requires at least a triple- basis set and triple excitations in the CC expansion87, 88, 89, 90. To investigate the perturbative triples correction in combination with the transcorrelated Hamiltonian, we evaluated the Be2 potential energy curve in the double- and triple- basis set. Furthermore, we compare results for the cc-pVDZ basis set with those obtained for the corresponding core-valence basis set cc-pCVDZ.
Fig. 2 shows that conventional CCSD results neither in the cc-pVDZ, nor in the cc-pCVDZ basis set yielded a qualitatively correct potential energy curve, because it is repulsive and no minimum could be identified. The inclusion of perturbative and full triples resulted in a stable dimer, but produces unphysical humps for inter nuclear distances slightly longer than the equilibrium bond length. The same holds for conventional CCSD in a triple- basis, whereas the inclusion of triple excitations improved potential energy curve.
The application of the transcorrelated Hamiltonian resulted in a well defined minimum already for the cc-pVDZ basis set in every CC model, but the curves are affected by a small unphysical humps. Transcorrelated CCSD(T) lowered the energy around the minimum and reduced the unphysical hump. The inclusion of full triple exications further improved the simulation accuracy.
As the basis set size increases, the difference between transcorrelated and conventional results decreases, since it converges to the conventional Hamiltonian in the complete basis set limit. Because of the size of the triple- basis, we were unable to carry out full-transcorrelated Hamiltonian calculations in that basis. As expected, transcorrelated CCSD still delivered the unphysical hump, whereas transcorrelated CCSD(T) and CCSDT produced a qualitatively correct potential energy curve.
Interestingly, in the core-valence double- basis set, the bonding energy is even further reduced, even though the basis set size is smaller than the triple- basis set. Accordingly, it has already been suggested to apply core-valence basis sets in the context of transcorrelation91, 57. The inclusion of nucleus-nucleus-electron correlation in the correlator may lift this requirement24, which however, leads to a more complicated correlator. It therefore appears to be in order to optimize specific core-valence atomic-orbital basis sets for the application in transcorrelated calculations, to support a rather simple and universal correlation factor.
The transcorrelated CCSDT potential energy curve coincides with the transcorrelated density matrix renormalization group minimum from Ref. 23, but the unphysical hump is reduced for CCSDT, indicating that the bond dimension was too small.
Furthermore, we investigated the captured correlation energy of CCSD and CCSD(T) with respect to CCSDT (see Fig. 3).
Fig. 3 shows that the electron correlation of CCSD is captured equally for the conventional and transcorrelated Hamiltonian in the double- basis. However, in contrast to conventional CCSD, the transcorrelated variant captured more correlation energy in the larger basis sets. By contrast, transcorrelated CCSD(T), however, recoverd less correlation energy as the conventional counterpart, similar to the observations that we made for the Be atom. The potential reason is the lifted symmetry of the two-body operator in combination with the approximation of the left cluster operator as . Hence, a bivariational optimization, as described in Ref. 62, could potentially yield a lower error in the correlation energy. Since the CC models presented in this study are all single reference, this also means that the main part of the electron correlation from the transcorrelated Hamiltonian is in the reference determinant. This also explains why the normal-ordering approximation works well, even in a dynamically correlated system such as Be2 (see Fig. 4).
Generally the error of the normal-order approximation is negligible, with respect to the huge memory requirements of the three-body operator.
5 Conclusions
We derived a perturbative triples approximation for the transcorrelated Hamiltonian, including three-body couplings. We showed that the required number of tensor contractions for CCSD, CCSD(T), CCSDT, and CCSDTQ for a Hamiltonian including a three-body operator drastically increases, in contrast to the conventional Hamiltonian. Code generation tools were required for an implementation to tackle this challenge.
Even for high-order CC the normal-ordering approximation, which hides the mean-field contributions of the three-body functions in the lower-body operators and omits the remaining functions, has a negligible effect on the energies. Hence, the number of required tensor contractions and, therefore, the computational scaling can be reduced to that of conventional CC. Based on the results for the Be atom and dimer, we can conclude that our transcorrelated CCSD(T) approximation improves transcorrelated CCSD, however to a smaller degree than for the conventional Hamiltonian. The trancorrelated Hamiltonian roughly increased the basis set convergence by two cardinal numbers on the -contraction scheme, as we showed for LiH dissociation and the fact that TC-CCSD(T) and TC-CCSDT have been able to produce a qualitatively correct potential energy curve for the dissociation of the Be2 dimer.
We also showed that simple universal correlation factors, as the one applied in this work, can result in an unbalanced treatment of the electron correlation, which may be cured by fitting proper atomic orbital basis sets, in view of the results obtained for core-valence polarization basis sets.
Even though CC is not a hermitian theory, a bivariational treatment of the amplitudes could even further increase the accuracy of the perturbative triples, since the de-excitation operator would match the amplitudes for the bra.
Acknowledgements
M.R. and M.M. gratefully acknowledge support by ETH Research Grant (no. ETH-43 20-2). This work was presented at the Faraday Discussion ”Correlated electronic structure” in London, UK, on 17 July.